# LLM-TabFlow: Synthetic Tabular Data Generation with Inter-column Logical Relationship Preservation

### v1

@@ 3.4 Compressed Data Generation
$\mathbf{H}_{t}=\mathbf{H}_{0}+\alpha(t)\boldsymbol{\eta},\quad\boldsymbol{\eta}\sim\mathcal{N}(0,\mathbf{I})$ (5)
@@
where $\nabla_{\mathbf{H}_{t}}\log p(\mathbf{H}_{t})$ is the learned score function, which estimates the gradient of the log-density and guides the denoising process. And $\alpha^{\prime}(t)$ represents the time derivative of the noise scaling function $\alpha(t)$, meaning it quantifies how the noise level changes over time. The term $\mathbf{W}_{t}$ denotes the standard Wiener process, introducing controlled stochasticity into the reverse dynamics. To achieve accurate reconstruction, the model is trained using denoising score matching. Specifically, a neural network $\boldsymbol{\eta}_{\phi}(\mathbf{H}_{t},t)$ is optimized to approximate the noise added during the forward process. The training objective is to minimize the following loss function (Karras et al., 2022):
$\mathcal{L}=\mathbb{E}_{\mathbf{H}_{0}\sim p(\mathbf{H}_{0})}\mathbb{E}_{t\sim p(t)}\mathbb{E}_{\boldsymbol{\eta}\sim\mathcal{N}(0,\mathbf{I})}\|\boldsymbol{\eta}_{\phi}(\mathbf{H}_{t},t)-\boldsymbol{\eta}\|^{2}_{2}.$ (7)
@@ 4 Experimental Settings
This section presents the experimental setup used to assess the performance of the proposed approach, as well as the state-of-the-art tabular data generation methods for comparison. It includes details on the evaluation datasets, the framework used for evaluation, and the hyperparameter configurations.
@@ 4.1 Datasets
Table 1: Statistics of the two evaluation datasets.
@@
Two real-world industrial datasets were adopted for experimental evaluation. The first dataset is an open-source dataset [^1] https://data.mendeley.com/datasets/8gx2fvg2k6/3 from the retail sector, denoted as Retail in this paper. The second dataset, denoted as Purchasing, is a proprietary dataset on procurement that comes from a manufacturing company. The statistics of these two datasets are presented in Table 1.
@@ 4.2 Baselines
We compare our proposed method, LLM-TabFlow, against five representative approaches, including:
@@
These methods span multiple generative paradigms, providing a diverse benchmark to evaluate how LLM-TabFlow performs relative to state-of-the-art techniques across different modeling approaches.
@@ 4.3.1 Data Fidelity
[…]
@@ 4.4 Experimental Setup
The two evaluation datasets were first split into training and testing sets, following the separation ratio specified in Table 1. Tabular data generation methods were then trained on the real training set, and the trained models were used to generate synthetic dataset with the same size as the real training set.
@@
In our experiments, we employed GPT-4 (Achiam et al., 2023) for column reasoning, enabling LLM-TabFlow to capture complex inter-column relationships effectively. To ensure a robust evaluation, we generated ten synthetic datasets using each trained tabular generation method. Each synthetic dataset was then used to train task-specific models—specifically, an XGBoost classifier in our experiments—for downstream tasks such as prediction. We then calculated the mean and standard deviation of the results.
@@
The hyperparameters of the prediction model—XGBoost classifier—were search within the following parameter space: number of trees: {100, 200, 300}; learning rate: {0.1, 0.01, 0.001}; maximum tree depth: {5, 10}; minimum child weight: {1, 10}; and gamma: {0.0, 0.5, 1.0}. Additionally, Grid Search was employed to identify the optimal hyperparameter combination for the XGBoost classifer.
Software and Hardware
@@
5 Experiments Results
@@
Figure 4: Radar Charts illustrating the performance of all methods on the two datasets across the six dimensions: Accuracy, Diversity, Utility, Privacy, Consistency, and Dependency. Values closer to the outermost circle indicate better performance. LLM-TabFlow achieves the overall best performance, with the most balanced results across all metrics.
This section presents the evaluation results for synthetic data produced by our proposed method—LLM-TabFlow—and baseline methods. Each evaluation is conducted ten times, repeating the process of generating and evaluating synthetic data. The results are reported as the results presented as the mean and standard deviation in the format ${\tt mean}\pm{\tt stddev}$.
@@ 5.1 Overall Evaluation
The overall performance of LLM-TabFlow and the five baseline methods on the two datasets is visualized as Radar charts, as illustrated in Figure 4. To provide a more comprehensive evaluation and clearer visualization, the inter-column relationship dimension is divided into two distinct aspects: Consistency and Dependency. As a result, the radar charts consist of six dimensions: Accuracy, Diversity, Utility, Privacy, Consistency, and Dependency. Surprisingly, as shown in Figure 4, SMOTE, originally proposed in early 2000s, still outperforms advanced deep generative models in several aspects such as accuracy, diversity, and inter-column relationships, while also achieving strong performance in utility on the Purchasing dataset. However, compared to deep generative models such as latent-space or PLM-based methods, this interpolation-based approach heavily relies on real input data for data sampling and often compromising data privacy. These limitations significantly restrict its applicability in industrial scenarios involving sensitive data. In contrast, advanced models such as TabDDPM and CTGAN perform slightly worse than SMOTE across almost all metrics. While TabSyn, a state-of-the-art method, surpasses SMOTE in privacy and utility, it struggles to fully capture consistency and dependency in the real data. Overall, models relying on latent representation exhibit significant limitations in preserving logical relationships. Notably, GReaT, a PLM-based model, excels in maintaining strong logical relationships but falls short in other critical dimensions, including data fidelity and utility. In contrast, our proposed approach—LLM-TabFlow—achieves near-perfect results across all six dimensions, ensuring consistency and dependency while excelling in fidelity, utility, and privacy. This demonstrates the robustness and versatility of LLM-TabFlow in generating high-quality realistic synthetic tabular data. We further examine the experimental results in each of the six dimensions in the following sections.
@@ 5.2 Accuracy and Diversity
Retailing Density Estimation 98.02±0.02 90.38±0.03 33.11±0.02 96.38±0.04 89.58±0.02 96.46±0.03
Pairwise Correlation 96.21±0.62 74.41±0.14 36.78±0.01 94.81±0.04 71.00±1.16 92.52±0.42
$\alpha$-Precision 93.54±0.01 88.90±0.14 0.00±0.00 98.48±0.10 82.18±0.15 99.15±0.12
Average Coverage Score 99.41±0.08 82.27±0.14 76.23±0.23 99.52±0.11 92.34±0.15 99.18±0.14
$\beta$-Recall 72.21±0.13 1.71±0.01 0.00±0.00 22.62±0.12 24.05±0.14 27.85±0.13
Purchasing Density Estimation 99.11±0.05 86.36±0.04 48.01±0.04 98.61±0.05 92.72±0.08 98.14±0.07
Pairwise Correlation 98.83±0.11 82.25±0.33 51.00±0.04 98.63±0.20 74.18±3.32 92.34±1.12
$\alpha$-Precision 98.01±0.32 73.67±0.22 4.16±0.01 96.73±0.37 88.00±0.36 97.29±0.25
@@
Figure 5: Density plots for the three continuous columns (item profit ratio, product price, and latitude), comparing the distribution of real data and their synthetic counterparts generated by different methods. Curves that more closely align with the real data indicate better performance. Both LLM-TabFlow and TabSyn exhibit distributions that closely match the real data, outperforming other methods.
@@
Figure 6: Distribution plots for the three categorical columns (shipping mode, payment type, and order status), comparing synthetic data to real data. Distributions that closely match the real data indicate superior performance. Both LLM-TabFlow and TabSyn exhibit distributions that are significantly closer to the real data compared to other methods.
@@
(f) Ours (LLM-TabFlow)
Figure 7: Heatmap illustrating the absolute divergence in pairwise column correlations between the synthetic and real data. Lighter colors indicate smaller differences and better alignment. SMOTE, TabSyn, and LLM-TabFlow exhibit the closest alignment with the real data, outperforming other methods.
@@
(f) Ours (LLM-TabFlow)
Figure 8: Visualization of the coverage score for both categorical and numerical columns using Radar charts. Labels starting with N and C represent numerical and continues columns around circles, respectively. LLM-TabFlow, TabSyn, and SMOTE significantly outperform the other three methods in terms of coverage across all evaluated columns.
@@
For data accuracy, we first utilize low-order statistics, such as column-wise density estimation and pairwise correlation scores (i.e., Pearson Correlation), to measure the general statistical properties of the columns. As shown in Table 2, SMOTE achieves the overall best performance on both datasets, followed closely by LLM-TabFlow and then TabSyn. Secondly, for high-order statistics, which capture more complex statistical relationships, the $\alpha$-Precision scores demonstrate that LLM-TabFlow outperforms all other methods, achieving 99.15% and 93.92% on the Retailing and Purchasing datasets, respectively. TabSyn and SMOTE follow closely, with values of 98.48% and 93.19%. To further examine the accuracy of synthetic data generated by various models, we analyzed the density distributions of representative columns in the Retailing dataset. First, we presneted the distribution of the three continuous columns—item profit ratio, product price, and latitude—in Figure 5. As shown in Figure 5, TabSyn and LLM-TabFlow perform the overall best, producing distributions that are nearly identical to the real data. While CTGAN and GReaT outperform than TabDDPM and generate distribution somewhat close to the real data, they exhibit noticeable gaps in peak regions, showing biases to the real data. SMOTE presents lower variations in these regions, but it is overall slightly less accurate than TabSyn and LLM-TabFlow. For categorical columns, such as shipping mode, payment type, and order status, as shown in Figure 6, LLM-TabFlow outperforms all other methods, accurately replicating category counts in these columns. TabSyn performs slightly worse than LLM-TabFlow but better than SMOTE. These results highlight LLM-TabFlow and TabSyn is superior to SMOTE for categorical data generation. Additionally, we visualize the absolute divergence in pairwise column correlations between the real Retailing dataset and its synthetic counterpart using heatmaps, providing a clear and intuitive representation of the differences. As presented in Figure 7, SMOTE, Tabyn, and LLM-TabFlow, though comparable to each other, achieve much better performance than other methods (e.g., CTGAN, TabDDPM, and GReaT) in terms of pairwise column correlation, with a closer match to the real data.
@@
For the data diversity, we continue by evaluating how well the synthetic data captures all modes present in the real data using the average coverage score described in Equation 9. The experimental results are presented in Table 2. As shown in Table 2, for single columns, LLM-TabFlow and SMOTE achieve the highest values on the Purchasing and Retailing datasets, respectively. In addition, we illustrate the coverage scores of synthetic data generated by each method for categorical and numerical columns in the Retailing dataset using radar charts. As illustrated in Figure 8, CTGAN excels in covering more modes for categorical data, while TabDDPM captures a broader range of numerical data. GReaT performs comparably for both numerical and categorical columns but fails to cover numerical column 7 (See 8(d)). LLM-TabFlow, Tabsyn, and SMOTE nearly cover all modes for both categorical and numerical columns. To assess the diversity of joint feature interactions, we used $\beta$-Recall scores. SMOTE, which explicitly balances minority classes, achieves the highest scores across both datasets, followed by LLM-TabFlow and GReaT, which rank second for each dataset, respectively. Notably, TabDDPM performs poorly on both accuracy and diversity metrics, indicating its inability to generate sufficiently diverse values. In summary, LLM-TabFlow strikes an excellent balance between accuracy and diversity, outperforming other models in most aspects, particularly in synthesizing continuous and categorical data.

### v2

@@ 3.4 Compressed Data Generation
$\mathbf{H}_{t}=\mathbf{H}_{0}+\alpha(t)\bm{\eta},\quad\bm{\eta}\sim\mathcal{N}(0,\mathbf{I})$ (9)
@@
where $\nabla_{\mathbf{H}_{t}}\log p(\mathbf{H}_{t})$ is the learned score function, which estimates the gradient of the log-density and guides the denoising process. And $\alpha^{\prime}(t)$ represents the time derivative of the noise scaling function $\alpha(t)$, meaning it quantifies how the noise level changes over time. The term $\mathbf{W}_{t}$ denotes the standard Wiener process, introducing controlled stochasticity into the reverse dynamics. To achieve accurate reconstruction, the model is trained using denoising score matching. Specifically, a neural network $\bm{\eta}_{\phi}(\mathbf{H}_{t},t)$ is optimized to approximate the noise added during the forward process. The training objective is to minimize the following loss function (Karras et al., 2022):
$\mathcal{L}=\mathbb{E}_{\mathbf{H}_{0}\sim p(\mathbf{H}_{0})}\mathbb{E}_{t\sim p(t)}\mathbb{E}_{\bm{\eta}\sim\mathcal{N}(0,\mathbf{I})}\|\bm{\eta}_{\phi}(\mathbf{H}_{t},t)-\bm{\eta}\|^{2}_{2}.$ (11)
@@ 4 Experimental Settings
This section outlines our comprehensive experimental framework for evaluating two key aspects: (1) LLM-based reasoning of inter-column relationships and (2) synthetic tabular data generation. We detail our methodology using real-world industrial datasets, benchmark comparisons against state-of-the-art generation techniques, and complete implementation specifications including hyperparameter configurations.
5 Experimental Settings
@@ 5.1 Datasets
Table 1: Statistics of the two evaluation datasets for generation tasks.
@@
Our evaluation framework assesses both logical reasoning generalizability and synthetic data generation quality. For reasoning tasks, we first fine-tune/prompt-design our model using two open-source datasets: Retail dataset (from DataCo(Wan, 2021) and MIMIC-III’s ICU stays data (Johnson et al., 2016b). We then evaluate on three test sets: (1) a private Purchasing dataset for procurement scenarios, (2) the UCI Adult dataset (Becker and Kohavi, 1996), and (3) the rest of whole MIMIC-III subsets(Johnson et al., 2016b) to validate the robustness of LLM-TabLogic. For synthetic data generation evaluation, we focus those two key supply chain datasets: the Retail dataset and the Purchasing dataset. The statistics of these two industrial datasets are presented in Table 1.
We utilize the Retail dataset (from DataCo(Wan, 2021) and MIMIC-III’s ICU stays data (Johnson et al., 2016b) for our analysis. Evaluation is performed on three test sets: (1) a proprietary Purchasing dataset for procurement scenarios, (2) the UCI Adult dataset (Becker and Kohavi, 1996), and (3) additional subsets of MIMIC-III. For synthetic data generation, we focus on two key supply chain datasets: the Retail dataset (Wan, 2021) and the Purchasing dataset. Statistics for these industrial datasets are summarized in Table 1.
@@ 5.2 Baselines
We compare our proposed method, LLM-TabLogic, against five representative approaches, including:
@@
These methods span multiple generative paradigms, providing a diverse benchmark to evaluate how LLM-TabLogic performs relative to state-of-the-art techniques across different modeling approaches.
@@ 5.3.1 Data Fidelity
where $X_{i}$ represents the values of feature $i$ in the real dataset, $X^{{}^{\prime}}_{i}(j)$ represents the $j$-th row for feature $i$, $\mathbb{1}{\left[\cdot\right]}$ is the indicator function that returns 1 if the condition inside the brackets is true and 0 otherwise, $N$ is the total number of synthetic data points.
@@ 5.4 Experimental Setup
For the LLM reasoning stage, only the meta data of each tables are input into LLM-TabLogic for logics inference. For the generation, the two evaluation datasets were first split into training and testing sets, following the separation ratio specified in Table 1. Tabular data generation methods were then trained on the real training set, and the trained models were used to generate synthetic dataset with the same size as the real training set.
@@
In our experiments, we evaluated the logical reasoning capabilities of both GPT-4 (Achiam et al., 2023) and DeepSeek-Va R1 (DeepSeek-AI et al., 2025) models for inferring inter-column relationships, leveraging their distinct strengths to capture complex tabular dependencies. First, we compared GPT-4 (Achiam et al., 2023) and DeepSeek-Va R1 (DeepSeek-AI et al., 2025) under deterministic conditions (temperature = 0.1), reporting averaged results from 10 independent responses to measure baseline performance. Second, to assess robustness, we systematically varied the temperature (0.1–0.9) for both models, analyzing how output stability scales with stochasticity. This dual approach not only highlights the comparative performance of the two LLMs but also quantifies their stability under varying levels of stochasticity. We evaluate LLM hallucination in logical relationship prediction using the F1 score across three benchmark datasets (Retailing, ICUstays from MIMIC-III, and Adult). Our methodology involves: (1) designing and validating prompts on the Retailing and ICUstays datasets, then (2) testing on the Adult, Purchasing and the rest of whole MIMIC-III datasets. The F1 score balances precision ($P=\frac{TP}{TP+FP}$) to quantify hallucination rates (penalizing false positives) and recall ($R=\frac{TP}{TP+FN}$) to assess coverage (penalizing false negatives), providing a unified measure of both prediction accuracy and completeness in inter-column logic inference.
For the data generation tasks, to ensure a robust evaluation, we generated ten synthetic datasets using each trained tabular generation method. And we report mean and standard deviation metrics from 10 sampling runs using the best-validated model. Each synthetic dataset was then used to train task-specific models—specifically, an XGBoost classifier in our experiments—for downstream tasks such as prediction.
@@
For the Data re
The hyperparameters of the prediction model—XGBoost classifier—were search within the following parameter space: number of trees: {100, 200, 300}; learning rate: {0.1, 0.01, 0.001}; maximum tree depth: {5, 10}; minimum child weight: {1, 10}; and gamma: {0.0, 0.5, 1.0}. Additionally, Grid Search was employed to identify the optimal hyperparameter combination for the XGBoost classifer. For the generative models, we finetune them following their hyperparameter settings from papers
Computational Resources
@@
6 Logical Reasoning Results
Figure 4: Performance comparison of GPT-4 and DeepSeek-Chat on tabular reasoning tasks. (a) shows F1 scores of GPT-4 and DeepSeek-Chat across industrial datasets. (b) presents the effect of temperature on F1 performance of both models on the Purchasing Datasets.
As shown in Table 4, our comparative analysis highlights consistent performance across datasets for both GPT-4 and DeepSeek-Chat. On the Retailing and Transfers datasets, both models achieve perfect F1 scores (1.00), which can be attributed to prompt designs that were specifically fine-tuned for these domains. In contrast, the remaining datasets—Purchasing, MIMIC-III (Avg), and Adult Income—represent unseen data not included during prompt tuning. On these more challenging and generalizable tasks, DeepSeek-Chat outperforms GPT-4 on both the Purchasing dataset (0.98 vs. 0.93) and MIMIC-III average (0.96 vs. 0.91), while both models again achieve perfect scores on Adult Income (1.00). These results underscore DeepSeek-Chat’s robust and adaptable reasoning performance across both fine-tuned and unseen tabular tasks.
Temperature Sensitivity Analysis
We investigate the impact of temperature variation ($\tau\in[0.1,1.0]$) on model performance by analyzing F1 score trends. DeepSeek-Chat demonstrates strong robustness, maintaining stable F1 scores across the full temperature range with a mean of $0.79\pm 0.02$. In contrast, GPT-4 exhibits a notable decline in performance, with its F1 score dropping from 0.82 at low temperatures to 0.71 at higher temperatures. This highlights DeepSeek-Chat’s greater stability in reasoning tasks under diverse sampling conditions, whereas GPT-4’s performance is more sensitive to decoding randomness.
7 Data Genertaoin Results
@@
Figure 5: Radar Charts illustrating the performance of all methods on the two datasets across the six dimensions: Accuracy, Diversity, Utility, Privacy, Consistency, and Dependency. Values closer to the outermost circle indicate better performance. LLM-TabLogic achieves the overall best performance, with the most balanced results across all metrics.
This section presents the evaluation results for synthetic data produced by our proposed method—LLM-TabLogic—and baseline methods. Each evaluation is conducted ten times, repeating the process of generating and evaluating synthetic data. The results are reported as the results presented as the mean and standard deviation in the format ${\tt mean}\pm{\tt stddev}$.
@@ 7.1 Overall Evaluation
The overall performance of LLM-TabLogic and the five baseline methods on the two datasets is visualized as Radar charts, as illustrated in Figure 5. To provide a more comprehensive evaluation and clearer visualization, the inter-column relationship dimension is divided into two distinct aspects: Consistency and Dependency. As a result, the radar charts consist of six dimensions: Accuracy, Diversity, Utility, Privacy, Consistency, and Dependency. Surprisingly, as shown in Figure 5, SMOTE, originally proposed in early 2000s, still outperforms advanced deep generative models in several aspects such as accuracy, diversity, and inter-column relationships, while also achieving strong performance in utility on the Purchasing dataset. However, compared to deep generative models such as latent-space or PLM-based methods, this interpolation-based approach heavily relies on real input data for data sampling and often compromising data privacy. These limitations significantly restrict its applicability in industrial scenarios involving sensitive data. In contrast, advanced models such as TabDDPM and CTGAN perform slightly worse than SMOTE across almost all metrics. While TabSyn, a state-of-the-art method, surpasses SMOTE in privacy and utility, it struggles to fully capture consistency and dependency in the real data. Overall, models relying on latent representation exhibit significant limitations in preserving logical relationships. Notably, GReaT, a PLM-based model, excels in maintaining strong logical relationships but falls short in other critical dimensions, including data fidelity and utility. In contrast, our proposed approach—LLM-TabLogic—achieves near-perfect results across all six dimensions, ensuring consistency and dependency while excelling in fidelity, utility, and privacy. This demonstrates the robustness and versatility of LLM-TabLogic in generating high-quality realistic synthetic tabular data. We further examine the experimental results in each of the six dimensions in the following sections.
@@ 7.2 Accuracy and Diversity
Retailing Density Estimation 98.02±0.22 90.38±0.23 33.11±0.12 96.38±0.24 89.58±0.32 96.46±0.23
Pairwise Correlation 96.21±0.62 74.41±0.24 36.78±0.31 94.81±0.44 71.00±1.36 92.52±0.42
$\alpha$-Precision 93.54±0.11 88.90±0.24 0.00±0.00 98.48±0.21 82.18±0.19 99.15±0.17
Average Coverage Score 99.41±0.18 82.27±0.34 76.23±0.23 99.52±0.31 92.34±0.45 99.18±0.24
$\beta$-Recall 72.21±0.13 1.71±0.21 0.00±0.00 22.62±0.32 24.05±0.21 27.85±0.23
Purchasing Density Estimation 99.11±0.15 86.36±0.34 48.01±0.44 98.61±0.35 92.72±0.38 98.14±0.47
Pairwise Correlation 98.83±0.21 82.25±0.33 51.00±0.34 98.63±0.20 74.18±3.22 92.34±1.12
$\alpha$-Precision 98.01±0.32 73.67±0.22 4.16±0.31 96.73±0.37 88.00±0.36 97.29±0.25
@@
Figure 6: Density plots for the three continuous columns (item profit ratio, product price, and latitude), comparing the distribution of real data and their synthetic counterparts generated by different methods. Curves that more closely align with the real data indicate better performance. Both LLM-TabLogic and TabSyn exhibit distributions that closely match the real data, outperforming other methods.
@@
Figure 7: Distribution plots for the three categorical columns (shipping mode, payment type, and order status), comparing synthetic data to real data. Distributions that closely match the real data indicate superior performance. Both LLM-TabLogic and TabSyn exhibit distributions that are significantly closer to the real data compared to other methods.
@@
(f) Ours (LLM-TabLogic)
Figure 8: Heatmap illustrating the absolute divergence in pairwise column correlations between the synthetic and real data. Lighter colors indicate smaller differences and better alignment. SMOTE, TabSyn, and LLM-TabLogic exhibit the closest alignment with the real data, outperforming other methods.
@@
(f) Ours (LLM-TabLogic)
Figure 9: Visualization of the coverage score for both categorical and numerical columns using Radar charts. Labels starting with N and C represent numerical and continues columns around circles, respectively. LLM-TabLogic, TabSyn, and SMOTE significantly outperform the other three methods in terms of coverage across all evaluated columns.
@@
For data accuracy, we first utilize low-order statistics, such as column-wise density estimation and pairwise correlation scores (i.e., Pearson Correlation), to measure the general statistical properties of the columns. As shown in Table 2, SMOTE achieves the overall best performance on both datasets, followed closely by LLM-TabLogic and then TabSyn. Secondly, for high-order statistics, which capture more complex statistical relationships, the $\alpha$-Precision scores demonstrate that LLM-TabLogic outperforms all other methods, achieving 99.15% and 93.92% on the Retailing and Purchasing datasets, respectively. TabSyn and SMOTE follow closely, with values of 98.48% and 93.19%. To further examine the accuracy of synthetic data generated by various models, we analyzed the density distributions of representative columns in the Retailing dataset. First, we presneted the distribution of the three continuous columns—item profit ratio, product price, and latitude—in Figure 6. As shown in Figure 6, TabSyn and LLM-TabLogic perform the overall best, producing distributions that are nearly identical to the real data. While CTGAN and GReaT outperform than TabDDPM and generate distribution somewhat close to the real data, they exhibit noticeable gaps in peak regions, showing biases to the real data. SMOTE presents lower variations in these regions, but it is overall slightly less accurate than TabSyn and LLM-TabLogic. For categorical columns, such as shipping mode, payment type, and order status, as shown in Figure 7, LLM-TabLogic outperforms all other methods, accurately replicating category counts in these columns. TabSyn performs slightly worse than LLM-TabLogic but better than SMOTE. These results highlight LLM-TabLogic and TabSyn is superior to SMOTE for categorical data generation. Additionally, we visualize the absolute divergence in pairwise column correlations between the real Retailing dataset and its synthetic counterpart using heatmaps, providing a clear and intuitive representation of the differences. As presented in Figure 8, SMOTE, Tabyn, and LLM-TabLogic, though comparable to each other, achieve much better performance than other methods (e.g., CTGAN, TabDDPM, and GReaT) in terms of pairwise column correlation, with a closer match to the real data.
@@
For the data diversity, we continue by evaluating how well the synthetic data captures all modes present in the real data using the average coverage score described in Equation 13. The experimental results are presented in Table 2. As shown in Table 2, for single columns, LLM-TabLogic and SMOTE achieve the highest values on the Purchasing and Retailing datasets, respectively. In addition, we illustrate the coverage scores of synthetic data generated by each method for categorical and numerical columns in the Retailing dataset using radar charts. As illustrated in Figure 9, CTGAN excels in covering more modes for categorical data, while TabDDPM captures a broader range of numerical data. GReaT performs comparably for both numerical and categorical columns but fails to cover numerical column 7 (See 9(d)). LLM-TabLogic, Tabsyn, and SMOTE nearly cover all modes for both categorical and numerical columns. To assess the diversity of joint feature interactions, we used $\beta$-Recall scores. SMOTE, which explicitly balances minority classes, achieves the highest scores across both datasets, followed by LLM-TabLogic and GReaT, which rank second for each dataset, respectively. Notably, TabDDPM performs poorly on both accuracy and diversity metrics, indicating its inability to generate sufficiently diverse values. In summary, LLM-TabLogic strikes an excellent balance between accuracy and diversity, outperforming other models in most aspects, particularly in synthesizing continuous and categorical data.